OpenStudy (anonymous):

The area of rectangle is 20x^2-27x-8 the length is 4x-1. What is the width?

OpenStudy (jdoe0001):

\(\bf length\cdot width=20x^2-27x-8\qquad length=4x-1\qquad thus \\ \quad \\ length\cdot (4x-1)=20x^2-27x-8\implies length=\cfrac{20x^2-27x-8}{4x-1}\)